Tom is driving up a very steep incline hill of about 20 degrees. Car has mass off 1200kg and a coefficient of friction of 0.4 with the road. At the start of the hill the car has a momentum of 30,000 kg•m/s. If car does 875,000J of work over the next 100m what is the new momentum of the car? (Reminder that the friction and Fp will create energy lost/negative work.)


1
Expert's answer
2021-05-03T10:29:10-0400

"A_2-A_1=\frac{p_2^2-p_1^2}{2m},"

"A_1=mg(\mu\cos\alpha+\sin\alpha)l,"

"p_2=\sqrt{2m(A_2-mg(\mu\cos\alpha+\sin\alpha)l+p_1^2}=31200~\frac{kg\cdot m}{s}."